OpenStudy (anonymous):

If cosecθ - sinθ = m and secθ - cosθ = n, then prove (m²n)2/3 + (mn²)2/3 = 1

OpenStudy (anonymous):

cosecθ - sinθ = m can be written as 1/ sinθ - sinθ = m (1- (sinθ)^2 )/ sinθ = m or (cosθ)^2 / sinθ = m similarly (sinθ)^2 )/ cosθ=n and so on...

OpenStudy (anonymous):

I'm having confusion with the raised to 2/3 part.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now find m^2* n = ((cosθ)^2 / sinθ )^2 * (sinθ)^2 )/ cosθ = (cosθ)^3

OpenStudy (anonymous):

similarly n^2*m = ((sinθ)^2 )/ cosθ )^2 *(cosθ)^2 / sinθ = (sinθ)^3 can you do now

OpenStudy (anonymous):

(m^2* n )^(2/3) = ((cosθ)^3 )^(2/3)=(cosθ)^2 and (n^2* m )^(2/3) = ((sinθ)^3 )^(2/3)=(sinθ)^2 so adding we have (m^2* n )^(2/3) + (n^2* m )^(2/3) = (cosθ)^2 + (sinθ)^2 =1

hartnn (hartnn):

 if that helps clear any confusion ^^